The cost of patient-days for a small town hospital is listed below. Answer the following questions.

Cost Formula: y = $0.894x + $2,931

  • i)  What is the variable cost? __________
  • ii)  What is the fixed cost? _________
  • iii)  What does x represent? ___________
  • iv)  What does y represent? ___________
  • v)  Whatistheexpectedcostfor47patient-days? (round to 2 decimal places) _________
  • vi)  How many patient-days were there if the total cost was $3,021.29? (round to the nearest day) ________




1
Expert's answer
2022-05-30T15:16:19-0400

Solution

i) Variable costs - costs, the value of which depends on sick days and their value increases with an increase in the number of sick days and is represented in the cost formula by a slope equal to $ 0.894

ii) Fixed costs - costs that do not depend on the size of sick days and with an increase or decrease in the number of sick days, their value remains unchanged and represents the free term in the cost formula equal to $ 2,931

iii) x is the number of sick days the patient spent in the hospital

iv) y - the total amount of costs incurred by the hospital for a certain number of sick days

v) instead of x, we substitute 47 in the cost formula and get:

0.894 * 47 + 2,931 = $2,973.02

vi) The total cost is $3,021.29 and we need to find the number of patient-days, which will be equal to:

3,021.29 = 0.894 * х + 2,931

90.29 = 0.894 * х

х = 101 patient-days